Circular No. 8241

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003jz, 2003ka, AND 2003kb Three apparent supernovae have been discovered: SN 2003jz by M.

Armstrong (cf. IAUC 8226), and SNe 2003ka and 2003kb by D. Singer, B. Beutler, and W. Li (via LOSS/KAIT; cf. IAUC 8236):

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003jz Nov. 13.18 9^{h}46^{m}28.^{s}67 +45^{o}45′04′′.7 17.5 1′′ E, 4′′ S
2003ka Nov. 16.2 23 03 43.28 +34 58 50.3 18.1 7′′.3 W, 7′′.9 S
2003kb Nov. 18.5  6 16 12.59 +57 03 12.9 18.0 8′′.4 W, 7′′.0 N

Additional approximate magnitudes from the respective discoverers (unless otherwise noted): SN 2003jz in UGC 5225, 2002 Nov. 30 UT, [19.0 (KAIT); 2003 Oct. 17.153, 16.0; Nov. 17.126, 17.5. SN 2003ka in MCG +06-50-20, Oct. 27.2 UT, [19.5; Nov. 18.2, 18.0. SN 2003kb in UGC 3432, Nov. 4.5, [19.5; 19.3, 18.0. Armstrong adds that nothing was visible at the location of SN 2003jz on Palomar Sky Survey red and blue plates.

SUPERNOVAE 2003jh AND 2003jg D. A. Howell, University of Toronto, reports that preliminary reduction

of a spectrogram (range 400–1000 nm), taken with the Clay 6.5-m Magellan Telescope (+ LDSS2) on Nov. 17.3 UT, shows that SN 2003jh (cf. IAUC 8236) is a type-IIn supernova. Hα has a narrow peak atop a broader base (FWHM 5000 km/s); the redshift measured from the narrow peak is z = 0.030, consistent with the redshift of the host, MCG –02-11-30 (z = 0.0296, as listed by NED). A spectrogram of SN 2003jg (cf. IAUC 8235) shows it to be a type-Ib/c supernova, several weeks past maximum light.

V838 MONOCEROTIS M. Tapia and P. Persi write: “Due to a transcription error in the

reduction process, the magnitudes of V838 Mon reported on IAUC 8235 are incorrect. The correct photometry, based on images obtained with the 2.1-m telescope and mid-infrared camera (CID) at San Pedro Martir on Oct. 5.5 UT is: [8.7 μm] = 1.37, [9.7 μm] = 0.99 [12.5 μm] = 0.56, [18.9 μm] = 0.01. The standard, β Gem, was observed through a similar airmass.”

2003 November 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green